OpenStudy (nilankshi):

five year hence,the age of jacob will be three times that of his son.five years ago.jacob's age was seven times that of his son.what is thair present age?

OpenStudy (sasogeek):

why did u post this in the history section? :o

OpenStudy (sasogeek):

Anyways, in 5 years, jacob's age = J+5 in 5 years, his son's age = S+5 but Jacob will be 3 times older than his son, therefore J+5=3(S+5) J+5=3S+15 J=3S+10 --------------(equation 1) 5 years ago, jacob's age = J-5 5 years ago, his son's age = S-5 but Jacob was 7 times older than his son, therefore J-5=7(S-5) J-5=7S-35 J=7S-30 --------------(equation 2) this means that 7S-30=3S+10 7S-3S=10+30 4S=40 S=10 this means that from equation 2, J=7S-30 J=7(10)-30 J=70-30 J=40 Hence their present ages are 70 and 10 :) if you're not clear about anything, do let me know :)